THIS IS FOR SUBSTITUTION EQUATION. ITS
X+3Y=-4
Y+X=0

Answer by funmath(2933)

THIS IS FOR SUBSTITUTION EQUATION. ITS
E1) X+3Y=-4
E2) Y+X=0
Solve E2 for y and substitute that into E1 and solve for x
y+x=0
y+x-x=0-x
y=-x
:
x+3(-x)=-4
x-3x=-4
-2x=-4
-2x/-2=-4/-2
x=2
Now substitute that into E2 and solve for y
y+2=0
y+2-2=0-2
y=-2
The solution: (x,y)=(2,-2)
Happy Calculating!!!